Can you please help me with this problem?
the cost of one serving of peanut butter (in cents) for a random sample of 19 jars of peanut butter was to be:
22 27 32 26 26 19 16
26 14 21 20 21 20 17
12 32 17 9 16
a) give the five-number summary ( low, Q1, median, Q3, high)
b) calculate the IQR..
Thank you for your help in advance
carol

Answer by stanbon(48568)

the cost of one serving of peanut butter (in cents) for a random sample of 19 jars of peanut butter was to be:
22 27 32 26 26 19 16
26 14 21 20 21 20 17
12 32 17 9 16
a) give the five-number summary ( low, Q1, median, Q3, high)
Ans: low = 9
Q1 = 16
median = 20
Q3 = 26
max = 32
---------------------------
b) calculate the IQR..
IQR = (Q3 - Q1) = 26 - 16 = 10
================
Cheers,
Stan H.